a train leaves the station at 1 pm. it travels at a rate of 72 mph. another train leaves the same station an hour later. it travels at rate of 90 mph the second train follows the same route as the first, on a track parallel to the first. how long will it take the secoong train to catch up to the first
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

a train leaves the station at 1 pm. it travels at a rate of 72 mph.
---
DATA:
rate = 72; time = x hrs. ; distance = 72x miles
---------------------------------------------------------
another train leaves the same station an hour later. it travels at rate of 90 mph the second train follows the same route as the first, on a track parallel to the first.
DATA:
rate = 90 mph ; time = x-1 hr. ; distance = 90(x-1) miles
---------------------
how long will it take the second train to catch up to the first?
-----
Equation:
distance = distance
72x = 90(x-1)
72x = 90x-90
18x = 90
x = 5 hrs
x-1 = 4 hrs (time it will take the 2nd train to catch the 1st train)
===================
Cheers,
Stan H.

Answer by mananth(16949)

first train speed = 72 mph
second train speed = 92 mph.
effective speed = 92-72 = 20 mph.
..
second train leaves 1 hour late.
by then first train has traveled 72 miles.
..
Time = distance / speed
=72/20
=3.6 hours.
second train catches up after 3.6 hours